OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

x^2-16 the first is binomial the second will be some trinomial

OpenStudy (anonymous):

\[(x-4)(x+4) = (x^{2}-16)\] Difference of squares! ->Binomial \[(x+4)(x+4) = (x+4)^{2}\] The whole square formula ->Trinomial

OpenStudy (anonymous):

x^2+8*x+16 x^2 -16 for the first one the last term the one without variable is same as second excep with a - sign. The middle term is non existant =0, and the first term the x^2 is same on both.

OpenStudy (anonymous):

I'm confused @timo86m

OpenStudy (anonymous):

x^2+8*x+16 = (x + 4)(x + 4) x^2 -16 = (x + 4)(x - 4) ^ same ^ in bottom one middle term is 0 on top it will be 8*x ^ last term in bottom one the number is same only the sign is -

OpenStudy (anonymous):

ohhh okahy! i understand what you're saying <3 thanks :)